# Order Parameter Discovery for Quantum Many-Body Systems

### v1

@@ Abstract
Quantum phase transitions offer profound insights into fundamental quantum phenomena and enhance our understanding of complex materials and systems. However, identifying quantum phase transitions in the absence of conventional order parameters poses a significant challenge. To address this, we utilize reduced fidelity susceptibility (RFS) vector field to construct phase diagrams of various quantum systems and then demonstrate its efficacy in reproducing the phase diagrams of established models with known order parameter. To this end, we propose a new method for discovering the necessary order parameters for a given quantum model and illustrate its capability by identifying a suitable order parameter for the Axial Next Nearest Neighbour Interaction (ANNNI) Model. Our analysis, which includes decomposing the observable into its eigen-projectors alongside the finite-size scaling, confirms that our method successfully can determine order parameters and thus its capable of characterizing quantum phase transitions
@@ I Introduction
Quantum phase transitions (QPTs) in many-body quantum systems pose considerable challenges for both theoretical modeling and experimental observation. Unlike classical phase transitions, which are driven by thermal fluctuations, QPTs are induced by changes in external parameters like magnetic fields or pressure, resulting in alterations to the ground state properties of a system.
@@
Various methods have been developed to detect and characterize both classical and quantum phase transitions. Among these is Landau-Ginzburg theory [1], designed to characterize phase transitions and critical phenomena using a field theory description. The formalism extends Landau’s mean-field theory by constructing a free energy functional and accounting for spatial variations of the order parameter, a quantity that reflects the broken symmetry of the low-temperature phase.
@@
While traditional theory effectively describes phase transitions, studying QPTs in finite-size systems requires more sophisticated tools. This need arises from the Kadanoff extended singularity theorem [2, 3], which highlights the relevance of non-analytic behavior in thermodynamic quantities near the critical point, even in finite-dimensional systems.
@@
Figure 1: Visualisation of the uses of the reduced fidelity susceptibility (RFS). Here we show it can be used for both Phase diagram Construction of quantum systems and order parameter discovery.
@@
To address this challenge, various methods such as the Renormalization Group (RG) [4, 5] and Finite-Size Scaling (FSS) theory [6] have been developed. RG operates by systematically removing less relevant degrees of freedom, highlighting the scale-independent features of a system. And thus, it provides insights into phase transitions by analyzing how physical systems behave across various length scales. This method involves coarse-graining the system and tracking the flow of coupling constants, with critical points emerging as fixed points in this flow. FSS also investigates how physical quantities vary with system size close to the critical point, facilitating the extraction of critical exponents from finite systems. Another notable approach involves employing machine learning techniques to detect phase transitions by identifying critical points and transition regions [7, 8, 9].
@@
An additional specialized tool for both the detection and characterization of phase transitions is the quantum state fidelity [10]. This concept, first applied to phase transitions in 1967 through the Anderson orthogonality catastrophe [11], measures the overlap between quantum states as system parameters change. Fidelity susceptibility [12, 13], a key concept in quantum information theory and quantum many-body physics, quantifies how sensitive a quantum state is to perturbations. It is particularly useful for detecting and characterizing quantum phase transitions, as it reveals sharp changes or singularities in fidelity near a quantum critical point, without relying on order parameters.
In this paper, we extend the use of the fidelity susceptibility from previous studies and construct a vector field using the reduced fidelity susceptibility (RFS). To this end, we characterize and detect different phases, specifically in the cluster Hamiltonian and Axial Next-Nearest Neighbor Ising (ANNNI) model [14, 15]. Not only does this approach allow us to demonstrate the proposed method, but also captures intricate details, such as the structure of ANNNI’s floating phase. A schematic picture of our method can be seen in Figure 1. Furthermore, we explore whether it is possible to identify an order parameter from a given Hamiltonian that accurately captures the phase transitions of a quantum system. Using the ANNNI model as an example, we provide a mathematical framework that enables the discovery of new order parameters for a model Hamiltonian, facilitating the understanding of different phases.
@@
Finally, we note that our method can provide efficient certification and verification processes for quantum simulations and phase transition detection [16] and circumvents the need for full tomography of the wave-function, relying instead on reduced density matrix to capture the relevant characteristics thermodynamic information.
@@
The outline of the paper is as follows. In Section II we outline the notation used in the paper and introduce the concept of fidelity and the RFS. This is followed by Section III, where we outline the mathematical framework, introduce the RFS vector field, and highlight its potential to be used to detect phase transitions. In Section IV, we highlight how the aforementioned mathematical framework can be used to determine an order parameter for a given quantum system. We showcase this in Section V by using the above framework to construct the phase diagram for both the ANNNI and cluster Hamiltonians and demonstrate the ability to discover new order parameters. We end with a discussion of applications of this framework in Section VI and potential use cases for future work.
@@ II Preliminary
Next, we introduce the concept of fidelity. Fidelity is a measure of the similarity between two quantum states and all its definitions embody the probability of distinguishing one quantum state from another.
@@
Following our discussion on the fidelity, we next consider a generic many-body parametric Hamiltonian $H(\lambda)=H_{0}+\lambda H_{I}$, with eigenvalue equation $H(\lambda)\ket{\psi_{k}\left(\lambda\right)}=E_{k}\ket{\psi_{k}\left(\lambda\right)}$, where the index $k=0$ identifies the ground state. In a perturbative approach w.r.t. a small parameter $\delta$, the Taylor expansion for the fidelity reads
@@
The expansion for the ground state of the perturbed Hermitian [^3] This result is commonly known as the Rayleigh–Schrödinger perturbation theory [19]. $H(\lambda)$ reads
@@
where $K_{k,0}\coloneqq\bra{\psi_{k}\left(\lambda\right)}H_{I}\ket{\psi_{0}\left(\lambda\right)}$. The latter leads to another well-known formulation for the susceptibility
@@ III Phase Diagram Construction
Building upon the previous section’s exploration of fidelity susceptibility, we introduce the concept of the reduced fidelity susceptibility (RFS) vector field. In doing so, we consider a general many-body Hamiltonian parameterized by the space $\mathcal{X}\subseteq\mathbb{R}^{2}$, of which its decomposition in terms of base and driving components reads
$H(\boldsymbol{\lambda})=$ $H_{0}+\lambda_{1}H_{1}+\lambda_{2}H_{2}$ (8)
with control parameters $\left(\lambda_{1}\quad\lambda_{2}\right)^{\top}\in\mathcal{X}$. We anticipate that the choice of the two-dimensional parameters space is convenient for the visual approach, but not fundamental. Let |$\ket{\psi_{0}\left(\boldsymbol{\lambda}\right)}\in\mathcal{H}$ denote the ground state of the Hamiltonian $H$ evaluated at $\boldsymbol{\lambda}\in\mathcal{X}$, with $\mathcal{H}$ denoting the underlying Hilbert space. The Hilbert space $\mathcal{H}$ is assumed bipartite, that is $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, for some subsystems $A$ and $B$. In addition, we assume the ground state is non-degenerate. Let $\rho_{0}(\boldsymbol{\lambda})$ denote the reduced density matrix (RDM) resulting from tracing out the subsystem $B$ for the ground state, so
@@
We call the terms $\partial^{2}f/\partial\delta_{k}^{2}$ the reduced fidelity susceptibility for the corresponding parameter $\lambda_{k}$. The adjective reduced is justified by the fact that we are considering RDMs instead of pure states as in (4). Similar forms of susceptibility have been considered in [22, 23]. As for the regular susceptibility, it is clear that the linear term in the perturbative Taylor expansion of $f$ vanishes for $\boldsymbol{\delta}=\mathbf{0}$.
We next obtain the vector field $P:\mathcal{X}\to\mathbb{R}^{2}$ (under sufficient smoothing assumptions, see details in subsection C.1) with rule
@@
$p(\boldsymbol{\lambda})=$ $\mathrm{Arg}\left(\mathbf{e}_{1}^{\top}P(\boldsymbol{\lambda})+\imath\mathbf{e}_{2}^{\top}P(\boldsymbol{\lambda})\right).$ (13)
@@
Figure 2: Visualisation of the fidelity vector (12) fields argument (a) and vector field (b) at a phase transition line. Here, we note that the argument drastically changes at the phase transition in (a) and the phase transition corresponds to a source in the vector field in (b).
The practical implementation of the construction is expanded in Appendix A while in subsection C.1 we argue on the differentiability of the fidelity.
@@ IV Order parameter discovery
$\underset{M\in\mathcal{S}_{m}}{\min}$ $-\frac{1}{|I^{+}|}\sum_{i\in I^{+}}\langle M\rangle_{i}^{2}+\frac{\gamma}{|I^{-}|}\sum_{j\in I^{-}}\langle M\rangle_{j}^{2},$
@@
with $\langle M\rangle_{i}\coloneqq\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})M)$ defining the expectation of $M$ at $\boldsymbol{\lambda}_{i}$, and $\gamma>0$ a tradeoff parameter. We denoted by $\mathcal{S}_{m}$ the set of Hermitian matrices of order $m$, which is also the order of the RDM $\rho_{0}(\boldsymbol{\lambda}_{i})$.
@@
Informally, the optimization favors the non-zero expectation $\langle M\rangle_{i}$ for the labels $y_{i}>0$ (ordered phase), whereas the quadratic term $\langle M\rangle_{i}^{2}$ penalizes non-zero expectations for the labels $y_{i}<0$ (disordered phase). In essence, this mechanism is mimicking the order/disorder behavior of the order parameters. The choice of $\gamma$ is sensible since extremal values $\gamma\gg 1$ or $0<\gamma\ll 1$ could render the problem degenerate (see Appendix B). The balance between the two expectation terms is regulated by the tradeoff parameter $\gamma>0$.
@@ V.1 The ANNNI model
In the phase diagram shown in Figure 3, three distinct phase transitions are observed. The first transition is an Ising-like transition between the ferromagnetic (FM) and paramagnetic (PM) phases. The second is a Kosterlitz-Thouless (KT) transition that occurs between the paramagnetic and floating phase (FP). The final transition is the Pokrovsky-Talapov (PT) transition, which separates the floating phase (FP) from the antiphase (AP).
For more information regarding each of these phases and the corresponding phase transitions and their ground states see Appendix F.
@@
Using the formalism as outlined in Section III, we first obtain the ground states of the Hamiltonian using the density matrix renormalization group (DMRG). For details on how the RDM was obtained or the DMRG algorithm, see Appendix D. Next, we consider a two-site RDM and calculate the RFS vector field and its corresponding angles. In Figure 4 (a), we plot the angle as given in (13) of the RFS vector field, as well as the theoretical Ising, KT, and PT phase transition lines, given in green, orange and blue respectively. We note there is a noticeable overlap between the theoretical phase transition lines and the angle values, with the angle of the vector fidelity abruptly changing at each phase transition. Comparing the theoretical phase transition lines to the vector field plotted in Figure 4 (b), we observe that a source in the vector field indeed corresponds to a phase transition.
Additionally, it is worth noting that the fidelity vector field surpasses previous simulations in its ability to capture intricate details that were previously unresolved in the floating phase [32]. Distinct scar-like features are present, with the sources of the vector field corresponding to each of these structures indicating the presence of the phase transition between the paramagnetic and floating phases, specifically the KT transition. Future work will aim to explore the structure and fully utilize the potential of the RDS vector field in analyzing such complex phenomena. In the subsequent section, we demonstrate how we apply this method to determine order parameters for a given quantum system.

### v3

@@ Abstract
Quantum phase transitions reveal deep insights into the behavior of many-body quantum systems, but identifying these transitions without well-defined order parameters remains a significant challenge. In this work, we introduce a novel approach to constructing phase diagrams using the vector field of the reduced fidelity susceptibility (RFS). This method maps quantum phases and formulates an optimization problem to discover observables corresponding to order parameters. We demonstrate the effectiveness of our approach by applying it to well-established models, including the Axial Next Nearest Neighbour Interaction (ANNNI) model, a cluster state model, and a chain of Rydberg atoms. By analyzing observable decompositions into eigen-projectors and finite-size scaling, our method successfully identifies order parameters and characterizes quantum phase transitions with high precision. Our results provide a powerful tool for exploring quantum phases in systems where conventional order parameters are not readily available.
@@ I Introduction
Quantum phase transitions (QPTs) in many-body quantum systems pose considerable challenges for theoretical modeling and experimental observation. Unlike classical phase transitions, which are driven by thermal fluctuations, QPTs are induced by changes in external parameters like magnetic fields or pressure, resulting in alterations to the ground state properties of a system.
@@
Various methods have been developed to detect and characterize classical and quantum phase transitions. Among these is Landau-Ginzburg theory [1], designed to describe phase transitions and critical phenomena using a field theory description. Its formalism extends Landau’s mean-field theory by constructing a free energy functional and accounting for spatial variations of the order parameter. The latter reflects the broken symmetry of the low-temperature phase.
@@
While the theory above effectively describes phase transitions, studying QPTs in finite-size systems requires more sophisticated tools. This need arises from the Kadanoff extended singularity theorem [2, 3], highlighting the relevance of non-analytic behavior in thermodynamic quantities near the critical point.
@@
Figure 1: Visualisation of the reduced fidelity susceptibility (RFS) uses. Here, we show that it can be used for both phase diagram construction of quantum systems and order parameter discovery.
@@
Various methods such as the Renormalization Group (RG) [4, 5] and Finite-Size Scaling (FSS) theory [6] have been developed to study QPTs in finite-size systems. RG operates by systematically removing less relevant degrees of freedom, highlighting the scale-independent features of a system. Thus, it provides insights into phase transitions by analyzing how physical systems behave across various length scales. This method involves coarse-graining the system and tracking the flow of coupling constants, with critical points emerging as fixed points of this flow.
FSS also investigates how physical quantities vary with the system size close to the critical point, facilitating the extraction of critical exponents from finite systems.
Yet another approach involves employing machine learning techniques to detect phase transitions by identifying critical points and transition regions [7, 8, 9, 10, 11].
@@
A specialized tool for detecting and characterizing phase transitions is the quantum state fidelity [12]. This concept, first applied to phase transitions in 1967 through the Anderson orthogonality catastrophe [13], measures the overlap between quantum states as system parameters change. Fidelity susceptibility [14, 15], a key concept in quantum information theory and quantum many-body physics, quantifies how sensitive a quantum state is to perturbations. It is particularly useful for detecting and characterizing quantum phase transitions, as it reveals sharp changes or singularities in fidelity near a quantum critical point, without relying on a design of tailored order parameters.
In this paper, we extend the use of the fidelity susceptibility from previous studies (e.g. [14, 15]) and construct a vector field using a reduced fidelity susceptibility (RFS). We use the RFS for the qualitative classification of the different phases. As we will show, in the RFS vector field, sources appear as phase transitions, while sinks are representative of the ground states for the corresponding phases. In addition, we devise an optimization problem, which identifies observables representing order parameters from the RFS vector field. The order parameter discovery gives a quantitative description of the critical point. A schematic picture of our method can be seen in Figure 1.
We demonstrate the efficacy of the proposed method on detecting QPT for three different models: (i) axial next-nearest neighbor Ising (ANNNI) model [16, 17] (due to its phase diagram), (ii) the cluster Hamiltonian model (due to its non-local description), and (iii) a physically motivated example given by a Rydberg atomic quantum simulator. In all the above models the proposed approach allows us to capture intricate details, such as the structure of ANNNI’s floating phase or the string order parameter in the cluster model. For (iii) we discover additional features that were not identified by prior studies: specifically, we observe ripple-like features at high values of the parameters swept, which might suggest the presence of a more complex phase, yet an open question previously discussed in [49, 51, 52].
@@
In summary, our mathematical framework establishes the phases diagram and enables the discovery of the order parameters for a Hamiltonian model, facilitating the understanding of different phases. We stress that our method can provide efficient certification and verification processes for quantum simulations and phase transition detection [18] and circumvents the need for full tomography of the wave function, relying instead on reduced density matrix to capture the relevant characteristic thermodynamic information.
@@
The outline of the paper is as follows. In Section II we outline the notation used in the paper and introduce the concept of fidelity and the RFS. This is followed by Section III, where we outline the mathematical framework, introduce the RFS vector field, and highlight its potential to be used to detect phase transitions. In Section IV, we highlight how the aforementioned mathematical framework can be used to determine an order parameter for a given quantum system. We showcase this in Section V by using the above framework to construct the phase diagram for the ANNNI model, the cluster Hamiltonian, and a one-dimensional Rydberg chain, and demonstrate the ability to discover new order parameters. We end with a discussion of applications of this framework in Section VI and potential use cases for future work.
@@ II Preliminary
Next, we introduce the concept of fidelity. Fidelity measures the similarity between two quantum states and all its definitions embody the probability of distinguishing one quantum state from another.
@@
Following our discussion on the fidelity, we next consider a generic many-body parametric Hamiltonian $H(\lambda)=H_{0}+\lambda H_{I}$, with eigenvalue equation $H(\lambda)\ket{\psi_{k}\left(\lambda\right)}=E_{k}(\lambda)\ket{\psi_{k}\left(\lambda\right)}$, where the index $k=0$ identifies the ground state. For clarity, we may omit the dependence of the various objects from the parameter $\lambda$, when it does not result in ambiguity. In a perturbative approach w.r.t. a small parameter $\delta$, the Taylor expansion for the fidelity reads
@@
The expansion for the ground state of the perturbed Hermitian [^3] This result is commonly known as the Rayleigh–Schrödinger perturbation theory [21]. $H(\lambda)$ is given as;
@@
where $K_{k,0}\coloneqq\bra{\psi_{k}\left(\lambda\right)}H_{I}\ket{\psi_{0}\left(\lambda\right)}$. This leads to another well-known formulation for the susceptibility
@@ III Phase Diagram Construction
Building upon the previous section’s exploration of fidelity susceptibility, we next introduce the reduced fidelity susceptibility (RFS) vector field concept. In doing so, we consider a general many-body Hamiltonian parameterized by the space $\mathcal{X}\subseteq\mathbb{R}^{2}$, of which its decomposition in terms of base and driving components reads
$H(\boldsymbol{\lambda})=$ $H_{0}+\lambda_{1}H_{1}+\lambda_{2}H_{2},$ (8)
with control parameters $\left(\lambda_{1}\quad\lambda_{2}\right)^{\top}\in\mathcal{X}$. We anticipate that choosing the two-dimensional parameters space is convenient for the visual approach, but not fundamental. Let |$\ket{\psi_{0}\left(\boldsymbol{\lambda}\right)}\in\mathcal{H}$ denote the ground state of the Hamiltonian $H$ evaluated at $\boldsymbol{\lambda}\in\mathcal{X}$, with $\mathcal{H}$ denoting the underlying Hilbert space. The Hilbert space $\mathcal{H}$ is assumed bipartite, that is $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, for some subsystems $A$ and $B$. In addition, we assume the ground state is non-degenerate. Let $\rho_{0}(\boldsymbol{\lambda})$ denote the reduced density matrix (RDM) resulting from tracing out the subsystem $B$ for the ground state, so
@@
We call the terms $\partial^{2}f/\partial\delta_{k}^{2}$ the reduced fidelity susceptibility for the corresponding parameter $\lambda_{k}$. The adjective reduced is justified by the fact that we are considering RDMs instead of pure states as in (4). Similar forms of susceptibility have been considered in [24, 25]. We also note that for the regular susceptibility, it is clear that the linear term in the perturbative Taylor expansion of $f$ vanishes for $\boldsymbol{\delta}=\mathbf{0}$. In Appendix J we expand on some key concepts in information geometry and its connection to the definition in (11).
We next obtain the vector field $P:\mathcal{X}\to\mathbb{R}^{2}$ (under sufficient smoothing assumptions) defined by;
@@
$p(\boldsymbol{\lambda})=$ $\mathrm{Arg}\left(\mathbf{e}_{1}^{\top}P(\boldsymbol{\lambda})+\imath\mathbf{e}_{2}^{\top}P(\boldsymbol{\lambda})\right),$ (13)
@@
Figure 2: Visualisation of the fidelity vector field (12) principal argument (a) and vector field (b) at a phase transition line. Here, we note that the argument drastically changes at the phase transition in (a) and the phase transition corresponds to a source in the vector field in (b).
The practical implementation of the construction is expanded in Appendix A, also, we argue on the patterns of the vector field in Appendix H.
@@ IV Order parameter discovery
$\underset{M\in\mathcal{S}_{m}}{\min}$ $\sum_{(i,j)\in I^{+}\times I^{-}}\left(-\frac{\langle M\rangle_{i}^{2}}{p_{i}}+\frac{\langle M\rangle_{j}^{2}}{p_{j}}\right),$
@@
with $\langle M\rangle_{i}\coloneqq\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})M)$ defining the expectation of $M$ at $\boldsymbol{\lambda}_{i}$, and $p_{i}\coloneqq\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})^{2})$ the purity of the same RDM. We denoted by $\mathcal{S}_{m}$ the set of Hermitian matrices of order $m$, which is also the order of the RDM $\rho_{0}(\boldsymbol{\lambda}_{i})$.
@@
Informally, the optimization favors the non-zero expectation $\langle M\rangle_{i}$ for the labels $y_{i}>0$ (ordered phase), whereas the quadratic term $\langle M\rangle_{i}^{2}$ penalizes non-zero expectations for the labels $y_{i}<0$ (disordered phase). In essence, this mechanism is mimicking the order/disorder behavior of the order parameters, for all pairs of opposing (w.r.t. phase) RDMs $(\rho_{0}(\boldsymbol{\lambda}_{i}),\rho_{0}(\boldsymbol{\lambda}_{j}))$, with $(i,j)\in I^{+}\times I^{-}$.
A special case of the problem in (IV) arises when we consider the reference density matrices $\rho_{+}$ and $\rho_{-}$, each representing a phase (i.e. ordered vs disordered). The formulation becomes
$\underset{M\in\mathcal{S}_{m}}{\min}$ $-\frac{\operatorname{Tr}\left(\rho_{+}M\right)^{2}}{\operatorname{Tr}\left(\rho_{+}^{2}\right)}+\frac{\operatorname{Tr}\left(\rho_{-}M\right)^{2}}{\operatorname{Tr}\left(\rho_{-}^{2}\right)},$
s.t. $\|M\|_{F}^{2}\leq 1.$ (16)
In Appendix C, it is shown that the objective of the latter is the quadratic form of an operator connected to the Gram-Schmidt process (GS) [31] – a method related to the $QR$ decomposition in linear algebra. For a density matrix $\rho$, we define $\widehat{\rho}\coloneqq\rho/\|\rho\|_{F}$ its associated norm-1 PSD. Also, for Hermitian matrices $A,B$ of the same order, we denote by $\langle A,B\rangle$ their Frobenius inner product. The special formulation leads to the close form solution
$M=$ $\frac{\widehat{\rho_{+}}-\left\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\right\rangle\widehat{\rho_{-}}}{\sqrt{1-\left\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\right\rangle^{2}}},$ (17)
when $\rho_{+}\neq\rho_{-}$. The solution is Hermitian as it is a real-linear combination of PSDs. One can immediately verify that $\operatorname{Tr}(\rho_{-}M)=0$, whereas
$\operatorname{Tr}(\rho_{+}M)=$ $\|\rho_{+}\|_{F}\sqrt{1-\left\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\right\rangle^{2}}>0,$ (18)
as expected. Geometrically, the solution in (17) can be interpreted as subtracting from $\widehat{\rho_{+}}$ its projection on $\widehat{\rho_{-}}$, hence the resulting Hermitian is orthogonal to $\rho_{-}$ and has a non-zero overlap with $\rho_{+}$ (when $\rho_{+}\neq\rho_{-}$). Moreover, the latter provides an understanding of the general problem. Specifically, the objective in (IV) corresponds to the sum of special objectives in (IV), for all pairs of opposing indices $(i,j)\in I^{+}\times I^{-}$.
@@ V.1 The ANNNI model
The phase diagram can be seen in Figure 3, where three phase transitions are present. The first is the Ising-like transition between the Ferromagnetic (FM) and Paramagnetic (PM) phases, the second is the Kosterlitz-Thouless (KT) transition between the paramagnetic and floating phase (FP), and the final is the Pokrovsky-Talapov (PT) phase transition between the floating phase and antiphase (AP).
For more information on each of these phases and the corresponding phase transitions and their ground states see Appendix G.
@@
Using the formalism as outlined in Section III, we first obtain the ground states of the Hamiltonian using the density matrix renormalization group (DMRG). Next, we consider a two-site RDM and calculate the RFS vector field and its corresponding angles. For details on how the RDM was obtained or the DMRG algorithm, see Appendix E. In Figure 4 (a), we plot the angle as given in (13) of the RFS vector field, as well as the theoretical Ising, KT, and PT phase transition lines, given in green, orange and blue respectively. We note there is a noticeable overlap between the theoretical phase transition lines and the angle values, with the angle of the vector fidelity abruptly changing at each phase transition. Comparing the theoretical phase transition lines to the vector field plotted in Figure 4 (b), we observe that a source in the vector field indeed corresponds to a phase transition. This source and sink-like behavior, as seen in Figure 4 (b) corresponds to local maxima and minima of Equation 11 and is further analyzed in Appendix H of the appendix.
Additionally, it is worth noting that the fidelity vector field surpasses previous simulations in its ability to capture intricate details that were previously unresolved in the floating phase [36]. Distinct ripples-like features are present, with the sources of the vector field corresponding to each of these structures indicating the presence of the phase transition between the paramagnetic and floating phases, specifically the KT transition. Future work will aim to explore the structure and fully utilize the potential of the RDS vector field in analyzing such complex phenomena. In the subsequent section, we demonstrate how we apply this method to determine order parameters for a given quantum system.